OpenStudy (anonymous):

find the derivative of the following...Using the chain rule f(x)=square root sin(ln(4x^3+1))

OpenStudy (anonymous):

\[f(x)=\sqrt{\sin(\ln(4x^3+1))}\]

OpenStudy (anonymous):

the derivative of \(\sqrt{f(x)}\) is \[\frac{1}{2\sqrt{f(x)}}\] so the first step is \[\frac{1}{2\sqrt{\sin(\ln(4x^2+1))}}\times \frac{d}{dx}[\sin(\ln(4x^2+1))]\]\]

OpenStudy (anonymous):

the first line i wrote was wrong it should be the derivative of \(\sqrt{f(x)}\) is \[\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

then the derivative of sine is cosine, so next step is \[\frac{1}{2\sqrt{\sin(\ln(4x^2+1))}}\times \cos(\ln(4x^2+1))\times\frac{d}{dx}[\ln(4x^2+1)\]

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

ok good two more steps

OpenStudy (anonymous):

ok I am stuck I am ghalf way through and now stuck

OpenStudy (anonymous):

ok we got the first part right?

OpenStudy (anonymous):

yes I understood that

OpenStudy (anonymous):

last job is \[\frac{d}{dx}[\ln(4x^2+1)]\]

OpenStudy (anonymous):

\[\frac{d}{dx}\ln(f(x))=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

so last step is \[\frac{8x}{4x^2+1}\]

OpenStudy (anonymous):

chain rule all the way down

OpenStudy (anonymous):

Ok then I got \[\frac{ -32x^2+8 }{ (4x^2+1)^2}\]

OpenStudy (anonymous):

hold the phone where did the numerator come from?

OpenStudy (anonymous):

you need the derivative of \(\ln(4x^2+1)\) that is all

OpenStudy (anonymous):

the derivative of \(\ln(4x^2+1)\) is \(\frac{8x}{4x^2+1}\) you do not need to take the derivative of that, it IS the derivative

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{\sin(\ln(4x^2+1))}}\times \cos(\ln(4x^2+1))\times\frac{d}{dx}[\ln(4x^2+1)\] last step is \[\frac{d}{dx}[\ln(4x^2+1)]=\frac{8x}{4x^2+1}\] and you are done

OpenStudy (anonymous):

ok so then the (8x)/4x^2+1) is the answer

OpenStudy (anonymous):

to that part, yes'

OpenStudy (anonymous):

entire answer is \[\frac{1}{2\sqrt{\sin(\ln(4x^2+1))}}\times \cos(\ln(4x^2+1))\times\frac{8x}{4x^2+1}\]

OpenStudy (anonymous):

oh gotcha why does the I was putting in the wrong order on my paper writing it out is why it wasn't making any sense sorry:)

OpenStudy (anonymous):

not too bad really, peal off the functions one at a time